OpenStudy (anonymous):

The sum of two numbers is 44, and the larger number is 2 more than the smaller number. What is the smaller number? If x = the smaller number and y = the larger number, then which of the following systems of equations represents the word problem? x + y = 44 and y = x + 2 x + y = 44 and y = 2x x + y = 44 and x = y + 2

OpenStudy (anonymous):

sum means add so x+y=44, the larger is 2 more than the smaller y = 2 + x y=x+2

OpenStudy (anonymous):

its the first one then

OpenStudy (anonymous):

yes